OpenStudy (anonymous):

Find the perimeter of a triangle with vertices (2,4), (-1,3) and (2,-3).

myininaya (myininaya):

This question has also been repeated. If you have questions about any of the work let me know.

OpenStudy (anonymous):

how did you get 3 radical 5 for one of the sides? i got 3 radical 10

myininaya (myininaya):

I used the distance formula for (1,-3) and (2,-3)

OpenStudy (anonymous):

the points are (-1,3) and (2,-3).

myininaya (myininaya):

sorry yah i meant (-1,3) it is hard looking back and forth

myininaya (myininaya):

\[\sqrt{(-1-2)^2+(3-(-3))^2}=\sqrt{(-3)^2+(6)^2}=\sqrt{9+36}=\sqrt{45} =\sqrt{9} \sqrt{5}\]

myininaya (myininaya):

\[=3 \sqrt{5}\]

OpenStudy (anonymous):

oh i did (3-(-3)=9, my bad. thank you.